OpenStudy (anonymous):

Find the slope of the line passing through the points (a,−a−8) and (a+h,−(a+7h)−8).

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

hartnn (hartnn):

here x1=a, y1=-a-8 x2=a+h, y2=-(a+7h)-8

OpenStudy (anonymous):

I did that but still having trouble getting the correct answer, thanks though

hartnn (hartnn):

y1-y2 = 7h x1-x2=-h slope = -7 what was the given correct answer?

OpenStudy (anonymous):

I wasn't making the h in the denominator negative. Thanks for your help!

hartnn (hartnn):

welcome ^_^